# A Competition Winning Deep Reinforcement Learning Agent in microRTS

### v1

@@ Abstract
1 Introduction
@@
Deep reinforcement learning (DRL) has proven to be powerful at solving complex problems requiring several steps to achieve a goal, such as Atari games (Mnih et al., 2013), continuous control tasks (Lillicrap et al., 2016), and even real-time strategy (RTS) games like StarCraft II (Vinyals et al., 2019). The StarCraft II grandmaster agent AlphaStar was trained with thousands of CPUs and GPUs/TPUs for several weeks. RTS games are particularly challenging for DRL for several reasons: (1) the observation and action spaces are large and varied with different terrain and unit types; (2) each unit type can have different actions and abilities; (3) each action can control several units at once; (4) rewards are sparse (win, loss, or tie) and delayed by possibly several thousand timesteps; (5) winning requires combining tactical (micro) and strategic (macro) decisions; (6) actions must be taken in real-time (i.e., the game won’t wait for the agent to take an action); (7) the agent might not have full visibility of the game state (i.e., fog of war); and (8) events in the game might be non-deterministic.
@@
microRTS (stylized as $\mu$RTS) is a minimalist, open-source, two-player, zero-sum RTS game testbed designed for research purposes (Ontañón, 2013). It includes many aspects of RTS games, simplified: different unit types, unit-specific actions, terrain, resource collection and utilization to build units, and unit-to-unit combat where units have different strengths and weaknesses. microRTS also supports fog of war and non-determinism; however, these were disabled for the IEEE-CoG 2023 microRTS competition.
@@
The IEEE microRTS competitions have been hosted at the Conference on Games (CoG) nearly every year since 2019 and at the Conference on Computational Intelligence and Games (CIG) before that since 2017 (Ontañón et al., 2018). Competitors submit an agent that plays against other submissions and baselines in a round-robin tournament on 12 different maps: 8 Open (known beforehand) and 4 Hidden (unknown until after the competition results are released). Agents are supposed to submit actions every step within 100 ms. Without GPU acceleration, this is a significant constraint for deep neural network agents.
@@
This paper describes how the RAISocketAI agent [^1] https://github.com/sgoodfriend/rl-algo-impls/tree/main/rl_algo_impls/microrts was trained and became the first DRL agent to win the microRTS competition by winning at CoG in 2023. The agent consists of 7 trained neural networks, taking 70 GPU-days to train. This significant training time combined with the general difficulty in debugging and fine-tuning a DRL implementation could explain why DRL hasn’t been competitive so far. We demonstrate that transfer learning to specific maps was critical to winning the competition. This strategy and our training framework can be a starting point for future research and competition agents.
@@
2 Related Work
2.1 MicroRTS-Py
Huang et al. (2021) released MicroRTS-Py [^2] https://github.com/Farama-Foundation/MicroRTS-Py, an OpenAI Gym wrapper for microRTS that includes a Proximal Policy Optimization (PPO) (Schulman et al., 2017) implementation trained on 1 of the Open maps (16x16basesWorkers). They added action composition, a shaped reward function, invalid action masking, IMPALA-CNN (a convolutional neural network with residual blocks), and trained against a diverse set of scripted agents. The agent achieved a 91% win rate on a single map against a diverse set of competition bots.
In their ablation studies, they found invalid action masking was essential to training an agent that could compete at the most basic level (82% win rate with invalid action masking, 0% without). Using the residual block network IMPALA-CNN architecture instead of the Atari Nature CNN by Mnih et al. (2013) got the win rate up the rest of the way.
They experimented with two different ways to issue player actions: Unit Action Simulation (UAS) and GridNet (Han et al., 2019). UAS calls the policy iteratively on each unit, simulating the game state after each unit action before issuing all unit actions combined to the game engine. GridNet computes the actions for all units in a single policy call by computing unit action logits for all grid positions and using a player action mask to ignore cells that don’t have any units owned by the player. UAS performed better than GridNet (91% vs 89%). Despite UAS’s better performance, MicroRTS-Py is deprecating UAS in favor of GridNet because of UAS’s more complex implementation and difficulty to incorporate self-play and imitation learning, both features important in RAISocketAI and our further work.
@@ 2.1 MicroRTS-Py
2.2 DeepMind AlphaStar
Vinyals et al. (2019)’s AlphaStar is a grandmaster-level AI trained with DRL to play the RTS game StarCraft II. They created an initial set of agents through imitation learning: supervised learning using a dataset of observations, actions, and rewards to train a policy to mimic actions from the dataset. The dataset was created by sampling replays of top-quartile human players. The supervised agents were rated in the top 16% of human players and used as starting points for DRL. They created a league-based framework to train multiple agents in parallel, each with different opponents to beat, thus creating a diverse set of training agents. AlphaStar was trained on 3072 TPU cores and 50,400 preemptible CPU cores for a duration of 44 days.
microRTS is a much simpler game than StarCraft II, both in game mechanics and simulation cost. Huang et al. (2021) trained on a single map for 300 million steps in less than 3 GPU-days. RAISocketAI trained on 10 maps for 1.5 billion steps in 70 GPU-days. While RAISocketAI didn’t use supervised learning to bootstrap the agent, our following work uses imitation learning to train a competitive agent.
@@ 2.3 Lux AI Kaggle competitions
The competition platform Kaggle hosts Simulation competitions where competitors submit agents that play against other submitted agents in a turn-based game environment. About once a year since 2020, Kaggle features an RTS-like Simulation competition: Halite, Lux AI, Kore, and Lux AI Season 2. Rules-based agents won Halite, Kore, and Lux AI Season 2. The first season Lux AI winning DRL agent by Pressman et al. (2021) has many similarities to MicroRTS-Py: GridNet action space, reward shaping, and an actor-critic training algorithm (IMPALA with additional UPGO and TD($\lambda$) loss terms, instead of PPO). Instead of training with a shaped reward function throughout training, Pressman et al. (2021) used shaped rewards on a smaller map before transitioning to sparse win-loss rewards on larger and competition-size maps. The top DRL agent by Limburg (2023) in the Lux AI Season 2 competition used a "DoubleCone" neural network backbone with critic and actor heads. DoubleCone is similar to ResNet’s backbone but the middle residual blocks are downscaled 4x to reduce inference time. RAISocketAI switches from shaped to sparse rewards during training and uses the DoubleCone architecture.
3 Methods
3.1 Neural network architecture
We reimplement much of MicroRTS-Py, including the PPO implementation, action composition, shaped reward function, invalid action masking, GridNet, self-play, and scripted bot training. RAISocketAI’s codebase [^4] https://github.com/sgoodfriend/rl-algo-impls supports multiple environments and reimplementing allowed the environment to fit into the existing codebase. We extended the observation representation in two ways: (1) walls and (2) unit destinations as invalid move targets in the invalid action mask. Only the unit’s current location was considered invalid in MicroRTS-Py; however, microRTS doesn’t allow units to move where another unit is moving into. Padded positions were represented as walls, which are impassable and noninteractive.
@@ 3.1 Neural network architecture
RAISocketAI loads 7 different policy networks. Only one network is used for a given map (Table 5). Networks are selected by (1) gathering all networks that are compatible with the map and its size, (2) prioritizing map-specific networks over size-specific networks, and (3) picking the highest priority network that can run within the allotted time on the current hardware. Multiple networks can be prioritized for the same map or map size. microRTS supports any map size (even non-square), and observations are padded to fit the policy network. Policy actions are clipped to fit the map size.
@@
RAISocketAI uses two different neural network backbones: Limburg (2023)’s DoubleCone(4, 6, 4) and a custom network (squnet). The actor head is a convolutional layer that outputs logits for unit actions at every position. A unit action is composed of independent discrete subactions: $D=\{a_{action\ type},$ $a_{move\ direction},$ $a_{harvest\ direction},$ $a_{return\ direction},$ $a_{produce\ direction},$ $a_{produce\ type},$ $a_{relative\ attack\ position}\}$. Invalid action masking sets logits to a very large negative number (thus zeroing probabilities and gradients) for actions that are illegal or would accomplish nothing (e.g., moving a unit to an occupied or reserved position) (Huang & Ontañón, 2020). This masking significantly reduces the action space per turn and makes training more efficient.
@@
While DoubleCone could support any map size, inference would likely exceed 100 milliseconds for larger maps. Therefore, for larger maps, we used a different architecture (squnet) which nests 3 downscaling blocks (Figure 4). This creates a network shaped like U-Net, but functionally similar to DoubleCone. This aggressive downscaling reduces the number of operations necessary during inference, especially for larger maps (Table 7).
@@
Instead of 1 value head, RAISocketAI uses 3 values heads for 3 different value functions: (1) shaped reward similar to MicroRTS-Py except each combat unit type is scaled by build-time (rewarding expensive units more), (2) win-loss sparse reward at game end (Tanh activation), and (3) in-game difference in units based on cost (similar to the reward function used by Winter (2021)). These 3 value heads are used to mix-and-match rewards over the course of training, generally starting with dense rewards using heads (1) and (3) and finishing with only win-loss sparse rewards by the end.
3.2 Base model training
We train using the PPO loss function from Schulman et al. (2017):
$L^{CLIP+VF+S}(\theta)$ $=\hat{\mathbb{E}}_{t}\left[L^{CLIP}(\theta)-c_{1}L^{VF}(\theta)+c_{2}S\left[\pi_{\theta}\right](s_{t})\right],$ (1)
@@ 3.2 Base model training
$L^{CLIP}(\theta)$ $=\min\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{old}}(a_{t}|s_{t})}\hat{A}_{t},\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{old}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right),$ (2)
$L^{VF}(\theta)$ $=\frac{1}{2}\left(V_{\theta}(s_{t})-\hat{V}_{t}\right)^{2},$ (3)
@@
where $c_{1}$ is the value loss coefficient, $c_{2}$ is the entropy coefficient, $S$ is an entropy bonus function, $\pi_{\theta}$ is the stochastic policy, $\pi_{\theta_{old}}$ is the rollout policy, $a_{t}$ is the action taken at time $t$, $s_{t}$ is the observation at time $t$, $\hat{A}_{t}$ is the advantage estimate, $\epsilon$ is the clipping coefficient, $V_{\theta}$ is the value function, $\hat{V}_{t}$ is the return estimate. We estimate the advantage using Generalized Advantage Estimation (GAE) (Schulman et al., 2016). We compute a separate advantage for each of the 3 rewards with independent $\gamma$ (discount factor) and $\lambda$ (exponential weight discount) for each reward. The 3 advantages are weighted summed together to get the final advantage estimate, which is used in computing the policy loss ($L^{CLIP}$).
@@
The reward weights, value loss coefficients (each value head has its own loss coefficient), entropy coefficient, and learning rate are varied on a schedule (Table 8). At the start of training, the policy loss heavily weighs the shaped reward advantage, and the value loss similarly weighs the shaped value head. By the end of training, both losses are weighted more towards the win-loss sparse reward. The entropy coefficient is also lowered at the end of training to discourage the agent from making random moves as the learning rate is lowered. The schedule specifies phases where these values are set and transitions between phases by changing values linearly based on timesteps.
@@
We first trained RAISocketAI with self-play both against itself and against prior versions of itself (Table 9). We reached a 91% win rate against the same bots MicroRTS-Py was benchmarked against. However, it only beat CoacAI (the 2020 competition winner) in 20% of games. The best performing agent of MicroRTS-Py nearly always beat CoacAI; however, the best versions using GridNet also usually lost against CoacAI. We fine-tuned the model through 3 iterations: (1) one-half of environments trained against CoacAI (Table 10); (2) one-half of environments trained against CoacAI or Mayari (2021 competition winner) split evenly and primarily trained on win-loss rewards (Table 11); and (3) same as before with action mask improvements and a GELU activation after the stride-4 convolution to match Limburg (2023)’s DoubleCone. By the end of fine-tuning, the model was winning 98% of games, including about 90% against each of CoacAI and Mayari.
3.3 Transfer learning
@@
The model so far had been trained on the 5 smaller Open maps. Using the fine-tuned model as a starting point, we trained additional models, each trained exclusively on 1 of 3 Open maps. NoWhereToRun9x8 is very different from the other maps with a wall of resources separating opponents. DoubleGame24x24 and BWDistantResources32x32 are larger than the maps the base model trained on. All 3 transfer learning runs used the same schedule: (1) warmup of sparse, win-loss reward weights linearly transitioned to a mixture of both shaped and sparse rewards, (2) middle phase of mixed rewards, (3) end phase of sparse rewards again at a lower learning rate (Table 12). We did an additional fine-tuning on NoWhereToRun9x8 using only the win-loss reward. These transfer learned agents exceeded 90% win rate on their respective maps, significant improvements over the base model (especially on BWDistantResources32x32 which started below 10%)
3.4 squnet training

### v2

@@ Abstract
Index Terms:
Machine learning, Games, Artificial Intelligence
publicationid: pubid: 979-8-3503-5067-8/24/\$31.00 ©2024 IEEE
I Introduction
@@
Deep reinforcement learning (DRL) has proven to be powerful at solving complex problems requiring several steps to achieve a goal, such as Atari games [1], continuous control tasks [2], and even real-time strategy (RTS) games like StarCraft II [3]. The StarCraft II grandmaster agent AlphaStar was trained with thousands of CPUs and GPUs/TPUs for several weeks. RTS games are particularly challenging for DRL for several reasons: (1) the observation and action spaces are large and varied with different terrain and unit types; (2) each unit type can have different actions and abilities; (3) each action can control several units at once; (4) rewards are sparse (win, loss, or tie) and delayed by possibly several thousand timesteps; (5) winning requires combining tactical (micro) and strategic (macro) decisions; (6) actions must be computed within a reasonable time window; (7) the agent might not have full visibility of the game state (i.e., fog of war); and (8) events in the game might be non-deterministic.
@@
microRTS (stylized as $\mu$RTS) is a minimalist, open-source, two-player, zero-sum RTS game testbed designed for research purposes [4]. It includes many aspects of RTS games, simplified: different unit types, unit-specific actions, terrain, resource collection and utilization to build units, and unit-to-unit combat where units have different strengths and weaknesses. microRTS also supports fog of war and non-determinism; however, these were disabled for the IEEE-CoG 2023 microRTS competition.
@@
The IEEE microRTS competitions have been hosted at the Conference on Games (CoG) nearly every year since 2019 and at the Conference on Computational Intelligence and Games (CIG) before that since 2017 [5]. Competitors submit an agent that plays against other submissions and baselines in a round-robin tournament on 12 different maps: 8 Open (known beforehand, Fig. 1) and 4 Hidden (unknown until after the competition results are released). Agents are supposed to submit actions every step within 100 ms. Without GPU acceleration, this is a significant constraint for deep neural network agents.
Refer to caption Refer to caption
basesWorkers8x8A FourBasesWorkers8x8
Refer to caption Refer to caption
NoWhereToRun9x8 basesWorkers16x16A
Refer to caption Refer to caption
TwoBasesBarracks16x16 DoubleGame24x24
Refer to caption Refer to caption
BWDistantResources32x32 (4)BloodBath.scmB
Figure 1: Open competition maps.
@@
This paper describes how the RAISocketAI agent [^1] https://github.com/sgoodfriend/rl-algo-impls/tree/main/rl_algo_impls/microrts was trained and became the first DRL agent to win the microRTS competition by winning at CoG in 2023. The agent chooses between 7 policy networks based on the map and compute capabilities. The significant training time (70 GPU-days) combined with the general difficulty in debugging and fine-tuning a DRL implementation could explain why DRL hasn’t been competitive so far. We demonstrate that transfer learning to specific maps was critical to winning the competition. This strategy and our training framework can be a starting point for future research and competition agents.
@@
II Related Work
II-A MicroRTS-Py
Reference [6] released MicroRTS-Py [^2] https://github.com/Farama-Foundation/MicroRTS-Py, an OpenAI Gym wrapper for microRTS that includes a Proximal Policy Optimization (PPO) [7] implementation trained on 1 of the Open maps (they used 16x16basesWorkers, which is the same as basesWorkers16x16A). They added action composition, a shaped reward function, invalid action masking, IMPALA-CNN [8] (a convolutional neural network with residual blocks), and trained against a diverse set of scripted agents. The agent achieved a 91% win rate on a single map against a diverse set of competition bots.
In their ablation studies, they found invalid action masking was essential to training a competitive agent (82% win rate with invalid action masking, 0% without). Using the residual block network IMPALA-CNN architecture instead of the Atari Nature CNN by [1] got the win rate up the rest of the way.
They experimented with two different ways to issue player actions: Unit Action Simulation (UAS) and GridNet [9]. UAS calls the policy iteratively on each unit, simulating the game state after each unit action before combining all actions to submit to the game engine. GridNet computes the actions for all units in a single policy call by computing unit action logits for all grid positions and using a player action mask to ignore cells that don’t have any units owned by the player. UAS performed better than GridNet (91% vs 89%). Despite UAS’s better performance, the MicroRTS-Py library is deprecating UAS in favor of GridNet because of UAS’s more complex implementation and difficulty to incorporate self-play and imitation learning, both features important in RAISocketAI and our further work.
@@ II-A MicroRTS-Py
II-B DeepMind’s AlphaStar
Reference [3]’s AlphaStar is a grandmaster-level AI trained with DRL to play the RTS game StarCraft II. They created an initial set of agents through imitation learning: supervised learning using a dataset of observations and actions to train a policy to mimic the actions from the dataset. The dataset was created by sampling replays of top-quartile human players. The supervised agents were rated in the top 16% of human players and used as starting points for DRL. They created a league-based framework to train multiple agents in parallel, each with different opponents to beat, thus creating a diverse set of training agents. AlphaStar was trained on 3072 TPU cores and 50,400 preemptible CPU cores for a duration of 44 days.
microRTS is a much simpler game than StarCraft II, both in game mechanics and simulation cost. Reference [6] trained on a single map for 300 million steps in less than 3 GPU-days. RAISocketAI trained on 10 maps for 1.5 billion steps in 70 GPU-days. While RAISocketAI didn’t use supervised learning to bootstrap the agent, our following work uses imitation learning to train a competitive agent.
@@ II-C Lux AI Kaggle Competitions
The competition platform Kaggle hosts Simulation competitions where competitors submit agents that play against other submitted agents in a turn-based game environment. Since 2020, Kaggle has featured 4 RTS-like Simulation competitions: Halite, Lux AI, Kore, and Lux AI Season 2. Rules-based agents won Halite, Kore, and Lux AI Season 2. The first season Lux AI winning DRL agent by [10] has many similarities to MicroRTS-Py: GridNet action space, reward shaping, and an actor-critic training algorithm (IMPALA with additional UPGO and TD($\lambda$) loss terms, instead of PPO). Instead of training with a shaped reward function throughout training, [10] used shaped rewards on a smaller map before transitioning to sparse win-loss rewards on larger and competition-size maps. The top DRL agent by [11] in the Lux AI Season 2 competition used a "DoubleCone" neural network backbone with critic and actor heads. DoubleCone is similar to ResNet’s backbone but the middle residual blocks are downscaled 4x to reduce inference time. RAISocketAI transitions from shaped to sparse rewards during training and uses the DoubleCone architecture.
III Methods
@@
Compared to [6], RAISocketAI’s biggest differences are (1) the DoubleCone neural network architecture, (2) adding self-play to the training regime, (3) using a training schedule that transitions from shaped to sparse rewards, (4) training on multiple maps, and (5) using transfer learning to train specialized models for specific maps. RAISocketAI loads 7 policy networks, but only uses one network at a time chosen based on map and compute capabilities (Table VI). Networks are selected by (1) gathering all networks that are compatible with the map and its size, (2) prioritizing map-specific networks over size-specific networks, and (3) picking the highest priority network that can run within the allotted time on the current hardware. microRTS supports any map size (even non-square), and observations are padded to fit the policy network. Policy actions are clipped to fit the map size.
We reimplement much of MicroRTS-Py, including the PPO implementation, action composition, shaped reward function, invalid action masking, GridNet, self-play, and scripted bot training. RAISocketAI’s codebase [^4] https://github.com/sgoodfriend/rl-algo-impls supports environments beyond microRTS and reimplementing allowed the environment to fit into the existing codebase. We extended the observation representation in two ways: (1) walls and (2) unit destinations as invalid move targets in the invalid action mask. Only units’ current locations were considered invalid move targets in MicroRTS-Py; however, microRTS doesn’t allow units to move where another unit is moving into. Padded positions were represented as walls, which are impassable and noninteractive.
III-A Neural Network Architecture
@@
RAISocketAI uses two different neural network backbones: [11]’s DoubleCone(4, 6, 4) (Fig. 2) and a custom network (squnet). The actor head is a convolutional layer that outputs logits for unit actions at every position. A unit action is composed of independent discrete subactions: $D=\{a_{\text{action type}},$ $a_{\text{move direction}},$ $a_{\text{harvest direction}},$ $a_{\text{return direction}},$ $a_{\text{produce direction}},$ $a_{\text{produce type}},$ $a_{\text{relative attack position}}\}$. Invalid action masking sets logits to a very large negative number (thus zeroing probabilities and gradients) for actions that are illegal or would accomplish nothing (e.g., moving a unit to an occupied or reserved position). This masking significantly reduces the action space per turn and makes training more efficient [12].
Figure 2: DoubleCone architecture.
@@
While DoubleCone could support any map size, inference would likely exceed 100 milliseconds for larger maps. Therefore, for larger maps, we used a different architecture (squnet) which nests 3 downscaling blocks. This creates a network shaped like U-Net, but functionally similar to DoubleCone. This aggressive downscaling reduces the number of operations necessary during inference, especially for larger maps where squnet does 8-fold fewer operations compared to DoubleCone on the largest 64x64 map.
@@
Instead of 1 value head, RAISocketAI uses 3 values heads for 3 different value functions: (1) shaped reward similar to MicroRTS-Py except each combat unit type is scaled by build-time (rewarding expensive units more), (2) win-loss sparse reward at game end (Tanh activation), and (3) in-game difference in units based on cost (similar to the reward function used by [13]). These 3 value heads are used to mix-and-match rewards over the course of training, generally weighing the dense reward heads (1 and 3) more heavily at the start of training and finishing with only the win-loss reward contributing at the end.
III-B Base Model Training
We train using the PPO loss function from [7]:
$L(\theta)$ $=\hat{\mathbb{E}}_{t}\left[L^{\text{CLIP}}(\theta)-c_{1}L^{\text{VF}}(\theta)+c_{2}S\left[\pi_{\theta}\right](s_{t})\right],$ (1)
$L^{\text{CLIP}}(\theta)$ $=\min\bigg{(}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{old}}(a_{t}|s_{t})}\hat{A}_{t},$
@@ III-B Base Model Training
$\qquad\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{old}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\right)\hat{A}_{t}\bigg{)},$ (2)
$L^{\text{VF}}(\theta)$ $=\frac{1}{2}\left(V_{\theta}(s_{t})-\hat{V}_{t}\right)^{2},$ (3)
@@
where $c_{1}$ is the value loss coefficient, $c_{2}$ is the entropy coefficient, $S$ is an entropy bonus function, $\pi_{\theta}$ is the stochastic policy, $\pi_{\theta_{old}}$ is the rollout policy, $a_{t}$ is the action taken at time $t$, $s_{t}$ is the observation at time $t$, $\hat{A}_{t}$ is the advantage estimate, $\epsilon$ is the clipping coefficient, $V_{\theta}$ is the value function, $\hat{V}_{t}$ is the return estimate. We estimate the advantage using Generalized Advantage Estimation (GAE) [14]. We compute a separate advantage for each of the 3 rewards with independent $\gamma$ (discount factor) and $\lambda$ (exponential weight discount) for each reward. The 3 advantages are weighted summed together to get the final advantage estimate, which is used in computing the policy loss ($L^{\text{CLIP}}$).
@@
The reward weights, value loss coefficients (each value head has its own loss coefficient), entropy coefficient, and learning rate are varied on a schedule. At the start of training, the policy loss heavily weighs the shaped reward advantage, and the value loss similarly weighs towards the shaped value head. By the end of training, both losses are weighted more towards the win-loss sparse reward. The entropy coefficient is also lowered at the end of training to discourage the agent from making as many random moves as the learning rate is lowered. The schedule specifies phases when these values are set and transitions between phases by changing values linearly based on timesteps.
@@
We first trained RAISocketAI with self-play both against itself and against prior versions of itself. We reached a 91% win rate against the same bots MicroRTS-Py was benchmarked against. However, it only beat CoacAI (the 2020 competition winner) in 20% of games. The best performing agent of MicroRTS-Py nearly always beat CoacAI; however, the best versions using GridNet also usually lost against CoacAI. We fine-tuned the model through 3 iterations: (1) one-half of environments trained against CoacAI; (2) one-half of environments trained against CoacAI or Mayari (2021 competition winner) split evenly and primarily trained on win-loss rewards; and (3) same as before with action mask improvements and a GELU activation after the stride-4 convolution to match [11]’s DoubleCone. By the end of fine-tuning, the model was winning 98% of games, including about 90% against each of CoacAI and Mayari.
III-C Transfer Learning
@@
Up to this point, the model had only been trained on the 5 smaller Open maps. Using the fine-tuned model parameters to jumpstart training [15], we trained additional models, each trained exclusively on 1 of 3 Open maps. NoWhereToRun9x8 is very different from the other smaller maps with a wall of resources separating opponents. DoubleGame24x24 and BWDistantResources32x32 are larger than the maps the base model trained on. All 3 transfer learning runs used the same schedule: (1) warmup of sparse, win-loss reward weights linearly transitioned to a mixture of both shaped and sparse rewards, (2) middle phase of mixed rewards, (3) end phase of sparse rewards at a lower learning rate. We did additional fine-tuning on NoWhereToRun9x8 using only the win-loss reward. These transfer learned agents exceeded 90% win rate on their respective maps, significant improvements over the base model (especially on, BWDistantResources32x32 which started below 10%).
III-D Squnet Training